A parking meter contains dimes and nickels worth ​$5.35 There are 65 coins in all. Find how many of each there are?
Found 2 solutions by ikleyn, richwmiller:
Answer by ikleyn(52786)

.
If you want to learn on how to solve coin problems, look into the lesson on this subject that were written specially for you:
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive
    - Three methods for solving standard (typical) coin word problems
    - More complicated coin problems
    - Solving coin problems mentally by grouping without using equations
    - Santa Claus helps solving coin problem
    - OVERVIEW of lessons on coin word problems


You will find there the lessons for all levels - from introductory to advanced,
and for all methods used - from one equation to two equations and even without equations.

Read them and become an expert in solution of coin problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Coin problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.



Answer by richwmiller(17219)

We will want an equation for the NUMBER of coins. d+n=65
We will also want an equation for the VALUE of coins. .10d+.05n=5.35
.10d+.05n=5.35
d+n=65
Let's get rid of decimals.
10d+5n=535
d+n=65
multiply by 5
5d+5n=65*5
10d+5n=535
5d+5n=325
subtract
5d=535-325=210
d=210/5=42
d+n=65
42+n=65
n=23
check
10d+5n=535
10*42+5*23=535
ok